Find the differential of y=(x+a)²


1
Expert's answer
2021-01-20T03:59:58-0500

Find the differential of y=(x+a)²

differential, dy = y' dx

Since "a" is a constant.

using chain rule to find y'

"{dy \over dx} = {dy \over du} × {du \over dx}"


Let u"=x + a"

then "{du \over dx} = 1"


y = u²

"{dy \over du} = 2u"


"\because" "{dy \over dx} = 2u . 1"


"dy =2(x +a)dx"